### v1

# A Quantitative Comparison of Pre-Trained Model Registries to Traditional Software Package Registries

Abstract.
Background: Software Package Registries (SPRs) are an integral part of the software supply chain. These collaborative platforms unite contributors, users, and packages, and they streamline package management. Much engineering work focuses on synthesizing packages from SPRs into a downstream project. Prior work has thoroughly characterized the SPRs associated with traditional software, such as NPM (JavaScript) and PyPI (Python). Pre-Trained Model (PTM) Registries are an emerging class of SPR of increasing importance, because they support the deep learning supply chain. Aims: A growing body of empirical research has examined PTM registries from various angles, such as vulnerabilities, reuse processes, and evolution. However, no existing research synthesizes them to provide a systematic understanding of the current knowledge. Furthermore, much of the existing research includes unsupported qualitative claims and lacks sufficient quantitative analysis. Our research aims to fill these gaps by providing a thorough knowledge synthesis and use it to inform further quantitative analysis.
Methods: To consolidate existing knowledge on PTM reuse, we first conduct a systematic literature review (SLR). We then observe that some of the claims are qualitative and lack quantitative evidence. We identify quantifiable metrics assoiated with those claims, and measure in order to substantiate these claims. Results: From our SLR, we identify 12 claims about PTM reuse on the HuggingFace platform, 4 of which lack quantitative validation. We successfully test 3 of these claims through a quantitative analysis, and directly compare one with traditional software. Our findings corroborate qualitative claims with quantitative measurements. Our two most notable findings are: (1) PTMs have a significantly higher turnover rate than traditional software, indicating a dynamic and rapidly evolving reuse environment within the PTM ecosystem; and (2) There is a strong correlation between documentation quality and PTM popularity. Conclusions: Our findings validate several qualitative research claims with concrete metrics, confirming prior qualitative and case study research. Our measures show further dynamics of PTM reuse, motivating further research infrastructure and new kinds of measurements.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
As the size and cost of developing deep learning (DL) models from scratch continue to rise, engineers are increasingly turning to adapt open-source Pre-trained Models (PTMs) as a cost-effective alternative (Jiang et al., 2022). PTM registries facilitate the reuse of open-source models by providing packages that include pre-trained weights, configuration, and documentation (Jiang et al., 2023b). Hugging Face has become a prominent PTM registry, comparable in popularity to traditional software registries like NPM and PyPI (Jiang et al., 2023b).
Prior research has empirically compared PTM registries to traditional software package registries such as NPM and PyPI, tackling diverse issues such as carbon emissions, model selection, and vulnerabilities (Jiang et al., 2023b; Castaño et al., 2023; Kathikar et al., 2023). Despite these efforts, no work has yet synthesized the existing knowledge of PTM registries. Furthermore, not all topics about PTM registries such as Hugging Face have been studied from both qualitative and quantitative perspectives. This lack of a comprehensive approach has led to a lack of quantitative validation, leaving numerous qualitative insights about PTM registries under-verified.
Figure 1. Overview of this work’s context and approach. Much is known, both qualitatively and quantitatively, about package reuse processes in traditional software package registries such as NPM and PyPI. Meanwhile, empirical data about the reuse of pre-trained deep neural network models (PTMs) is emerging. This work provides the first systematic literature review on PTM reuse, focused on extracting the claims present in the prior work (RQ1) and providing quantitative evaluation of the un-quantified and under-quantified claims (RQ2).
Our work aims to bridge the knowledge gap concerning the synthesis of extensive research on PTM reuse facilitated by Hugging Face, and provide further measurements to validate the prior qualitative and quantitative claims. An overview is shown in Figure 1. Our work consists of two parts: First, we conduct a systematic literature review to transform qualitative insights about PTM package reuse on Hugging Face into quantitative metrics. Second, we evaluate the robustness of these insights through quantitative validation. This approach enhances our understanding of package reuse dynamics on Hugging Face compared to traditional SPRs.
Our findings indicate that most of the claims from prior qualitative results are correct. Of the 3 prior claims we quantitatively evaluated, 2 were supported by our results, and 1 was supported by 1 measurement and not supported by the other. Our measurement of library usage indicated a preference for using the Transformers library when creating descendents of models, with 80% of descendents of models choosing to utilize Transformers. Our measurement of package turnover indicates that HuggingFace has a significantly higher turnover rate than traditional package registries. Our measurement of model popularity and descendent count found a correlation between the two, further supporting claims that popularity is a driver of model selection. Finally, we found a strong correlation between documentation quality and PTM popularity, with the top 1000 models outperforming the bottom 1000 in documentation.
Our contributions are:
  • We conduct a systematic literature review on PTM reuse in the Hugging Face registry, and extract a list of qualitative claims from prior work. (section 4)
  • We map the qualitative claims to quantitative measurements. We use these measurements to validate the prior findings, via comparison of our quantitative measurements of Hugging Face to representative traditional SPRs. (section 5)
  • Our work provides recommendations for future work on developing tools to analyze and keeping datasets updated to support further investigations on the PTM supply chain. (section 7)
Significance for Software Engineering: Prior work has made both quantitative and qualitative claims about software engineers’ reuse of PTMs. We systematically synthesized this knowledge through a literature review and developed quantifiable metrics to corroborate the qualitative claims. Our findings confirm several qualitative results about PTM reuse, and also quantify the dynamic reuse environment within the PTM ecosystem. Our results will inform research infrastructure and the development of new metrics to guide and refine PTM development and reuse.
Table 1. Example metrics used to characterize traditional software package registries. The metrics in this table are used as guidance when developing metrics to measure the claims in Table 3.
Metric Description & Implication Package Registries Example Works
User Reach The amount of ecosystem controlled by a fraction of its maintainers. NPM, PyPI, Cargo, Elm, CRAN (Ahmed et al., 2017; Bommarito and Bommarito, 2019; Liu et al., 2022; Blanthorn et al., 2019; Zhang et al., 2020; Gu et al., 2023; Zimmermann et al., 2019)
Dependency Degree The number of dependencies among software packages. NPM, PyPi, Cargo, CPAN, CRAN, NuGet, Packageist, RubyGems, Maven (Zerouali et al., 2019a; Pinckney et al., 2023; Wittern et al., 2016; Zerouali et al., 2018; Bommarito and Bommarito, 2019; Decan et al., 2019; Liu et al., 2022; Blanthorn et al., 2019; Gu et al., 2023; Zerouali et al., 2019b; Decan et al., 2018; Zimmermann et al., 2019; Stringer et al., 2020; Wu et al., 2023; Kula et al., 2014)
Popularity The frequency of use for top packages, indicating the concentration of usage. NPM, PyPI, OpenStack (Wittern et al., 2016; Dey and Mockus, 2018; Zerouali et al., 2019b)
Technical Lag The delay in adopting new updates. NPM, PyPI, Maven, Cargo (Zerouali et al., 2019a; Pinckney et al., 2023; Wittern et al., 2016; Zerouali et al., 2018; Decan et al., 2018; Stringer et al., 2020)
2. Background and Related Work
In this section, we first explain how software package registries facilitate the reuse of software (section 2.1). We then discuss the rise of Deep Neural Networks (DNNs) and their presence in software engineering reuse as PTM packages (section 2.2). We also detail how software package registries have been measured in previous works, and the limited effort towards quantitatively evaluating PTM registries (section 2.3). Our work aims to advance the state of the art when it comes to quantitative evaluation of PTM registries.
2.1. Software Package Registries
Figure 2. The Software Supply Chain. Engineers contribute Packages to model registries, Packages evolve through internal development and external dependence on other Packages, and are used by downstream Reusers who incorporate them into Systems.
Figure 2 depicts the software supply chain. Software Package Registries (SPRs) serve as collaborative hubs that connect package contributors, reusers, and the packages themselves, facilitating software reuse. These registries are important for engineers because they provide comprehensive information that significantly enhances downstream software development.
SPRs act as platforms for creators to upload and share their software packages, featuring version control to allow users to access previous package versions as needed (npm, [n. d.]; pyp, [n. d.]). These platforms promote package discoverability and perceived quality through user engagement tools like comments, likes, and ratings. Additionally, collaboration is fostered by discussion threads and version management, enhancing user interaction and visibility. Registries also provide comprehensive metadata, improving package visibility in search results and aiding in package selection. Tools for package downloading, bundling, and managing versions and updates streamline the reuse process, simplifying package lifecycle management.
Recently, PTM package registries (e.g., Hugging Face (Wolf et al., 2020), PyTorch Hub (pyt, [n. d.]), ONNX Model Zoo (onn, [n. d.])) emerged to support efficient development of AI systems (Davis et al., 2023; Jiang et al., 2022). PTM packages include traditional components (e.g., documentation, dependencies). However, they also include additional DNN-specific components, such as pre-trained weights, training dataset, and model architecture (Jiang et al., 2023b). Figure 3 depicts the structural similarity in package reuse between traditional and PTM packages.
Figure 3. Reuse processes of traditional software and of PTMs, as reported by Jiang et al. (Jiang et al., 2023b). Reuse processes are similar, suggesting that SPR measurements and trends may be similar.
2.2. Deep Neural Networks and Pre-Trained Model Packages
Deep Neural Networks (DNNs), comprising numerous hidden layers, have gained popularity as a cutting-edge solution for complex problems in various fields, such as image recognition in autonomous vehicles (Garcia et al., 2020) and AI voice assistant systems (Nasirian et al., 2017). Developing and training these models from scratch requires significant time and resources (Rae et al., 2021; Hoffmann et al., 2022; Patterson, 2022). For example, a Llama-2-70B model needs 1720K GPU hours to train from scratch (Touvron et al., 2023). To overcome this, engineers increasingly opt for using Pre-Trained Models (PTMs), which allows them to forego the lengthy and resource-intensive initial training phase. By leveraging PTMs, they can focus on a much shorter training period to fine-tune the models for specific tasks (Jiang et al., 2023b; Davis et al., 2023; Jiang et al., 2023a).
Within the context of PTMs, the term “software reuse” refers to the reuse of models along with their training configurations and weights, i.e., reuse of the PTM packages. Figure 3 shows the comparison of reuse process between PTM and traditional packages. Reusing DNNs as PTM packages mirrors traditional practices of software package reuse, where existing software components are integrated and adapted rather than built from scratch (Jiang et al., 2022; Gong et al., 2023a). This practice can enhance efficiency and reducing development time (Gopalakrishna et al., 2023). This adaptation involves not only the models but also their pre-trained states, which can be adjusted for specific applications, thereby constituting a combination of model reuse and customization rather than traditional software package reuse alone.
2.3. Measurements of Software Package Registries
Research on software package registries has traditionally focused on quantifying various aspects (Zerouali et al., 2019a; Decan et al., 2018; German et al., 2010). However, PTM registry research has primarily focused on qualitative and small-scale quantitative measurements.
2.3.1. General Comparisons:
The measurement of software package registries is an established research area within traditional software. Table 1 shows the metrics used in prior work e.g., user reach (Zimmermann et al., 2019), license (German et al., 2010), and technical lag (Decan et al., 2018). These studies provide critical insights for software engineers, aiding in effective package selection and reuse, and for researchers, deepening understanding of software supply chains and registry dynamics.
This research approach has been adapted to PTM registries recently (Jiang et al., 2024a, 2023b). Early studies have explored how PTMs are reused and adapted, examining both qualitative and quantitative aspects. For instance, research has started to analyze contribution patterns and the reuse dynamics specific to PTM registries, such as those found in TensorFlow Hub and Hugging Face (Jiang et al., 2023b; Miao et al., 2017; Davis et al., 2023). However, there are still gaps in understanding the evolution and reuse patterns in the PTM registries (Jiang et al., 2023b, 2024b). Our work bridges this gap by providing additional quantitative measurements and comparing our results to traditional SPRs.
2.3.2. To Explain or Quantify Phenomena:
Despite these efforts, there remains a gap in research that connects qualitative observations with quantitative measurements across these registries.
Recent studies within traditional SPRs typically start with qualitative observations that are later supported by quantitative data. Issues such as package obsolescence and the spread of vulnerabilities, initially observed quantitatively, have been rigorously quantified in ecosystems like NPM (Mujahid et al., 2023; Abdalkareem et al., 2017). These investigations form a crucial basis for comprehending how features of software registries impact the practices of software development.
In the research domain of PTM registries, similar investigations are needed. For example, Jiang et al. provide a comprehensive analysis on the risks while using PTMs, and the PTM reuse process (Jiang et al., 2022, 2023b). As a follow-up work, they also collected both qualitative and quantitative data on PTM naming practices (Jiang et al., 2024a). This line of research is crucial for understanding PTM package registries. However, their qualitative insights have not been substantiated by quantitative analysis. We address this gap by conducting a systematic literature review and deriving quantifiable metrics to assess the claims from previous studies.
3. Knowledge Gap and Research Questions
To summarize the knowledge gap, we lack a cohesive understanding on PTM package reuse that synthesizes prior research on Hugging Face. Additionally, there is a lack of some quantitative measurements, which reduce our understanding of the registry.
To address the gap of knowledge synthesis, we ask:
  • RQ1
    What claims about package reuse on Hugging Face are made by prior research?
By gathering and analyzing these claims, we aim to convert the qualitative data into a set of quantifiable metrics. This effort will not only enrich our understanding of Hugging Face as a PTM platform but also enhance the comparability of data across different software ecosystems. Specifically, this analysis enables future researchers to establish metrics that facilitate the comparison of Hugging Face and traditional software registries similarities and differences in package reuse practices.
Answering RQ1 also prepares us for further empirical inquiry:
  • RQ2
    Do the qualitative claims about package reuse (PTM) on Hugging Face hold up when quantified?
3.1. Overview of Methodology
Figure 1 shows the overview of our work’s context and approach. We answered these questions through three steps:
  1. (1)
    We conducted a systematic literature review (SLR) on PTM reuse, aiming to compile a comprehensive list of claims about package reuse.
  2. (2)
    We synthesized these claims into quantifiable measurements, performed the measurements, and then compared them to the previous findings.
  3. (3)
    Where possible, we compared these to those from traditional software to see differences between the reuse patterns in different registries.
The detailed method per RQ is in the corresponding sections.
4. RQ1: What claims about package reuse on Hugging Face are made by prior research?
Finding 1. The Systematic Literature Review (SLR) identified 12 quantifiable claims about package reuse on Hugging Face. Finding 2. The claims are distributed among five quantifiable categories of methods: small- and large-scale quantitative measurements, qualitative surveys, interviews, and case studies. Finding 3. After this classification, four quantitatively unevaluated claims remained, within the categories of design trends, documentation and understanding, and selection considerations. This finding shows a gap that we address in RQ2.
To enhance our understanding of the existing claims related to PTM reuse within package registries, we followed empirical standards and conducted a systematic literature review (Ralph et al., 2020; Ralph and Baltes, 2022; Borenstien et al., 2009; Keele et al., 2007). Our initial step involved a pilot study to define the scope of our review (section 4.1.1). A systematic literature review entails five steps: identification of research, study selection, study quality assessment, data extraction, and data analysis (Keele et al., 2007; Moher et al., 2010). Figure 4 illustrates the results of each step. We detail them next.
Figure 4. Systematic literature review process. Step 1 details the search query used. The subsequent filtering and distinct claims, along with the summarized categories, are discussed in section 4.2.
4.1. Methods
4.1.1. Pilot Study
We define the scope of our review by conducting a pilot study. In the pilot study, we first search for papers about “pre-trained model reuse” using Google Scholar and looked at the first three results, which were (Jiang et al., 2022, 2023b; Gong et al., 2023a). The papers indicated that Hugging Face is the only “open” model registry (Jiang et al., 2022) and is the most popular model registry. Additionally, Hugging Face hosts the largest number of PTM packages, and provides useful tools to facilitate PTM reuse. We then decided to scope down our study on Hugging Face model registry specifically to represent the PTM supply chain, as indicated by Jiang et al. (Jiang et al., 2024b, 2023b).
4.1.2. Search Strategy and Query
The goal of our search is to identify papers that are relevant to the categories of PTM reuse, the PTM supply chain, or the PTM ecosystem. Informed by our pilot study ( section 4.1.1), the final search query we used is indicated in Figure 4. These search queries gave us 45 papers. We then removed duplicate entries, reducing the number of papers to 31.
To verify the efficacy of our search queries, we employed papers from the pilot study as benchmarks to assess each query’s retrieval effectiveness. We ensured that all papers identified in the pilot study were also retrieved by our final search query. This step was essential to confirm the robustness of our search strategy, ensuring it was capable of capturing the most relevant studies. Such a comprehensive approach allowed for an exhaustive review of the literature concerning PTM reuse within its operational ecosystem.
4.1.3. Selection Criteria
The goal of our selection criteria is to identify the most relevant and rigorously supported research that specifically addresses the context and impact of PTM reuse. During our study, we applied two types of criteria: inclusive and exclusive.
Inclusion Criteria We applied the following inclusion criterion: a paper is included if it describes the reuse of models within a specific PTM registry. This criterion excluded papers that apply PTMs to specific tasks, reducing our set from 31 papers to 20.
Exclusion Criteria Our exclusion criterion was that we excluded non-primary sources. We exclude works whose claims are not substantiated directly through qualitative or quantitative methods. This reduced our set from 20 papers to 18.
Table 2. The detailed definition and examples of each claim category we extracted from the literature review.
Claim Category Definition Examples
Motivation claims Articulates the rationale behind a paper’s problem statement, illustrating why the work is significant and worthy of investigation. The reuse of pre-trained models introduces large costs and additional problems of checking whether arbitrary pre-trained models are suitable for the task-specific reuse or not.” (Gong et al., 2023a)
With the commodification of AI, and NLP in particular…[we] need easy to use, no-code tools for understanding AI artifacts.” (Piktus et al., 2023)
Work claims Refers to the assertions a paper makes based on its collected data and analyses. Hugging Face’s popularity has exponentially increased over time, which is evident from the upward trends in the number of new models, likes, commits, unique authors, and discussions aggregated monthly.” (Castaño et al., 2023)
Engineers follow specific naming practices and encounter challenges that are specific to PTM naming.” (Jiang et al., 2024a)
4.1.4. Data Extraction
Once we identified the most relevant and rigorously supported research that specifically addresses the context and impact of PTM reuse, the next step is to extract “claims” from these papers that provide evidence of qualitative or quantitative methods that could inform our quantitative study. A “claim” in this context refers to a statement or assertion made in a research paper, which is supported by evidence.
The data extraction involve four steps. First, two co-authors went through a paper from the pilot study together to get an agreement of the data extraction process. The goal of this process was to identify and extract all claims that might be tangentially related to PTM reuse. Second, they individually extracted the claims from the papers. This involved reading each paper to identify the key claims, with an emphasis placed on the abstract, introduction, and, if available, finding boxes of each paper. Exact quotations from the papers were extracted. Third, the two co-authors met and presented their extracted claims from each paper, with an explanation of why it was chosen, and a discussion of the relevance of the claim if it was not immediately obvious. A total of 256 claims were discussed in this step. Finally, for each paper, they discussed which claims were the most descriptive of PTM reuse and discarded the rest. This selection step resulted in a total of 49 claims.
4.2. Analysis and Results
We categorized these claims into two categories: (1) “Motivation claims” and (2) “Work claims”. The detailed definitions and examples of each claim are shown in Table 2. This classification process yielded 19 motivation claims and 30 work claims. One of our goals (RQ2) was to substantiate prior findings with further quantitative measurements, so we chose to focus more deeply on the work claims. Within this subset, we identified overlapping claims and consolidated them, resulting in a refined set of 12 distinct claims. These consolidated claims are detailed in Table 3.
The primary aim of our SLR is to summarize the existing claims about package reuse on Hugging Face and to extract quantifiable measurements from these claims. From the consolidated set of 12 claims, we categorized the basis of each claim into one of five methods: small- and large-scale quantitative measurements, qualitative surveys, interviews, and case studies. Small-scale measurements involved less than 10% of a population, whereas large-scale measurements encompassed more than 10%. After this classification, five quantitatively unevaluated claims remained: one concerning design trends, two regarding selection considerations, and two about documentation and understanding.
The categorization result and extracted themes are detailed in Table 3. Our results provide a thorough knowledge synthesis which are then used to answer RQ2 (section 5).
Table 3. Consolidated themes and claims collected from our systematic literature review. Small-Scale measurements refer to measurements made on a selection of a population less than 10% of the overall size, Large-scale measurements are measurements made on a selection of a population that is more than 10%, Survey refers to a survey, Interviews refer to interviews, Case study refers to either an examination of a specific case or the creation of a model to verify the claim. $\dagger$ : The claim basis is not a large-scale quantification.
Themes Claims Claim Basis Works
Trends in Design A small group of contributors owns popular models. Large-Scale Measurement (Castaño et al., 2023)
The Transformers library increases the accessibility of PTM creation and downstream reuse. Large-Scale Measurement, Case Study (Kathikar et al., 2023; Wolf et al., 2020; Jiang et al., 2024b; Castaño et al., 2023)
$\dagger$ (1) The Transformers library improves the process of PTM evolution. Case Study (Wolf et al., 2020)
Forking repositories introduces low severity vulnerabilities. Large-Scale Measurement (Kathikar et al., 2023)
Selection Considerations $\dagger$ (4) DL-specific attributes such as model architecture, performance, reproducibility, and portability affect PTM selection and reuse. Interviews (Jiang et al., 2023b)
$\dagger$ (2) The traditional attribute of Popularity affects PTM selection and reuse more than Maintenance and Quality. Interviews (Jiang et al., 2023b)
Repository Lifecycle and Maintenance Models receive perfective maintenance over time, with high-maintenance models tending to be more popular, larger, and better documented. Large-Scale Measurement (Castaño et al., 2023)
Documentation and Understanding Model properties are under-documented across Hugging Face. Small-Scale and Large-Scale Measurement, Survey (Kathikar et al., 2023; Taraghi et al., 2024; Montes et al., 2022; Gong et al., 2023a)
$\dagger$ (3) Documentation quality impacts model selection. Survey, Case Study (Taraghi et al., 2024; Montes et al., 2022; Castaño et al., 2024)
Naming models is inconsistent and can inadequately represent model architectures. Large-Scale Measurement, Survey (Jiang et al., 2024a)
Dataset documentation is correlated with dataset popularity. Small-Scale Measurement (Yang et al., 2024)
Downstream Usage Hugging Face is an exponentially growing platform. Large-Scale Measurement (Castaño et al., 2023)
5. RQ2: Do the qualitative claims about package reuse (PTM) on Hugging Face hold up when quantified?
Finding 4. Our study shows that the Transformers library is preferred in over 80% of PTM descendants, surpassing PyTorch. The rise of the SafeTensors library underscores a shift toward prioritizing security in PTM development.
Finding 5. Figure 6 reveals that Hugging Face has a significantly higher package turnover rate than traditional software registries indicative of a fast-paced, innovation-driven PTM ecosystem. Finding 6. There is a correlation between model popularity and descendant count, indicating that while popular models have more descendants, other factors influence model selection decisions. Finding 7. There is a strong correlation between documentation quality and PTM popularity, with the top 1000 models significantly outperforming the bottom 1000 in documentation, highlighting its importance in model selection.
To answer this question, we first derive quantifiable metrics from the claims we extracted from our SLR (section 4) on Hugging Face (section 5.1). Then we present the available datasets and the specific data we used from each (section 5.2). Subsequently, we present our methods and results for measurement on metrics for each claims (section 5.3section 5.5).
5.1. Metrics Developed from Claims
In this section, we explain how we developed quantifiable metrics from the claims identified in our SLR (section 4). The metrics were specifically designed to quantify the hypothesized results inferred from these claims. Drawing on traditional software engineering practices, we adopted metrics that have been previously used to evaluate similar claims in other contexts if available.
Particular attention was paid to metrics that are widely recognized and have been frequently cited in the literature, as well as those that have been implemented across various traditional software registries. This approach ensures that our metrics are grounded in established methodologies and are robust enough to provide meaningful insights.
Table 4 shows the relationship between the claims extracted and the corresponding metrics, along with the expected measurements. This mapping both validaties the claims and contextualizes their implications within the context of package reuse on Hugging Face.
5.2. Available Datasets
The section presents the PTM package datasets (section 5.2.1) and traditional software package dataset (section 5.2.2) we used in our work.
5.2.1. PTM Datasets
In the PTM literature, there are four datasets available publicly: HF Model Metadata (Van Strien, 2022), PTMTorrent (Jiang et al., 2023c), HFCommunity (Ait et al., 2023), and PeaTMOSS (Jiang et al., 2024b).
  1. (1)
    HF Model Metadata provides a snapshot of 10,406 HuggingFace model metadata as of 11/2022, including details of model label, README and length of each README file (Van Strien, 2022).
  2. (2)
    PTMTorrent encompasses a snapshot of five model hubs, totaling 15,913 PTM packages as of 08/2023, all formatted in a uniform data schema to facilitate cross-hub mining (Jiang et al., 2023c).
  3. (3)
    HFCommunity is an offline relational database constructed from the data at the Hugging Face Hub. It allows for queries on the repositories hosted within the Hugging Face platform (Ait et al., 2023).
  4. (4)
    PeaTMOSS offers comprehensive metadata on 281,638 PTM packages as of 10/2023, including 281,276 from Hugging Face and 362 from PyTorch Hub, along with details on 28,575 GitHub projects that use PTMs as dependencies and 44,337 links from these GitHub repositories back to the PTMs they depend on (Jiang et al., 2024b).
In this work, we primarily used the PeaTMOSS dataset, as the accessibility of metadata allowed for easier measurements. We also used the HF Model Metadata and PTMTorrent datasets for longitudinal trends in the turnover metric (section 5.4), and an April 2024 recent snapshot of the most popular models from the Hugging Face Hub API for the same reason. The detailed data used for each measurement are presented in their respective sections.
5.2.2. Traditional package datasets
To directly compare measurements between the PTM registry and traditional Software Package Registries (SPRs), we utilized the Ecosyste.ms software package dataset for traditional packages, following the approach outlined in prior work (Schorlemmer et al., 2024). This dataset provides a set of free and open resources for those working to sustain and secure open source software. Ecosyste.ms publishes open data and APIs that map software interdependencies, as well as providing data on the usage, creation, and potential impact of packages.
In this work, we used the version of the dataset from October 2023, as this is the closest in time to when the PeaTMOSS dataset was created. This includes the usage of data from two software package registries: PyPI and NPM.
5.3. $C_{1}$: The Transformers library increases the accessibility of PTM creation and downstream reuse.
Table 4. This table displays the relationships between Qualitative Claims, the Metric(s) used to evaluate them, and the Hypothesized Results. Where applicable, a reference is made to the traditional software prior work and metric that informed our metric. Note that not every metric contains a reference to a traditional software prior work that uses this metric as no analog exists. For those metrics, we utilize the Goal-Question-Metric process to develop metric(s) associated with the claim. Since the goal of these measurements is to substantiate existing claims from prior work, we include a hypothesis about what the quantitative measurement will be if the claim is true.
Qualitative Claim Metric Hypothesized Result
$C_{1}$: The Transformers library improves the process of PTM evolution. Preservation rate of Libraries to descendents. The preservation of Transformers as a library to its descendents will be greater than that of other libaries.
$C_{2}$: The traditional attribute of Popularity affects PTM selection and reuse more than Maintenance and Quality. Turnover of top Packages over time (Dey and Mockus, 2018). Models with high popularity remain popular over time (“rich get richer”).
Descendent amount of models Models with high popularity have a larger number of descendent models.
$C_{3}$: Documentation quality impacts model selection. Popularity of PTMs based on their documentation quality. Models with more information are more discoverable and therefore more popular (Zerouali et al., 2018).
$C_{4}$: DL-specific attributes such as model architecture, performance, reproducibility, and portability affect PTM selection and reuse. Popularity of PTMs based on their Attributes (Zerouali et al., 2018). Popular architectures, high performance, description of where the PTM came from, ease of use, and size all impact model popularity.
5.3.1. Method
We present the metric we developed for claim 1. Metric 1: Preservation rate of libraries to descendents: Some models support multiple libraries. If the claim holds, then descendants of those models should make use of (and thus support) the more reuse-friendly libraries. We consider each library $L$ in turn to assess how frequently its descendants continue to use it. The ingredients of our measure are: the library $L$ being assessed, the set of base models $B$ that use library $L$ and at least one other library, the set $D_{b}$ of direct descendants of base model $b\in B$, and a function $S(d,L)$ that returns 1 if descendant $d\in D_{b}$ supports library $L$, otherwise 0. We can then calculate the preservation rate $P_{L}$ of a library $L$ as:
(1) $P_{L}=\frac{\sum_{b\in B}\sum_{d\in D_{b}}S(d,L)}{\sum_{b\in B}|D_{b}|}$
The numerator is the total descendants supporting library $L$, calculated as: $\sum_{b\in B}\sum_{d\in D_{b}}S(d,L)$. The denominator is the total count of direct descendants across all base models in set $B$: $\sum_{b\in B}|D_{b}|$.
This equation will give the percentage of direct descendants that support library $L$. The raw count of descendants supporting the library can be found using the numerator.
5.3.2. Results
Metric 1: Preservation rate of libraries to descendents: As depicted in Figure 5, the Transformers library has the highest survival rate from a parent model to its descendant among all libraries used. Specifically, over 80% of descendant models continue to employ the Transformers library, establishing it as the preferred choice for generating descendant models. Transformers is also the most popular library (has the largest population of models that use it). Contrarily, despite previous studies suggesting high prevalence, PyTorch was not favored among PTM descendants, a departure from claims within our systematic literature review that both Transformers and PyTorch are dominant on Hugging Face (Castaño et al., 2023).
Figure 5. The usage proportion of the top-10 libraries that PTMs utilizing at least two different libraries use on Hugging Face. For PTM packages that leverage support at least two libraries, most packages support the transformers library, followed by the Hugging Face promoted SafeTensors library. Most other libraries have little usage in comparison. In contrast to previous work (Castaño et al., 2023), PyTorch is not one of the most popular library to be supported when a PTM package supports more than one library.
Instead, our findings highlight the Transformers and SafeTensors libraries as the most prevalent, suggesting a community shift towards prioritizing security, as SafeTensors is designed to replace the commonly used Python pickle library with a more secure container for deploying models. We attribute this discrepancy to prior studies’ broader approach, which did not differentiate between model types and analyzed only a single snapshot of Hugging Face, thus lacking the detailed analysis presented here. Our results suggest that PTM synthesizers are willing to compromise on functionality and portability to ensure the distribution of more secure PTM packages.
5.4. $C_{2}$: Popularity Affects PTM Selection and Reuse More than Other Trad’l. Attributes
Our claim interpretation is that popular models are more likely to be used, so that the “rich get richer”. We examined this claim with two metrics. First, we measure the stability (i.e., non-turnover) of the top PTM packages over time, expecting it to be low (popular packages are used directly). Second, we measure the correlation between popularity and the number of descendants of top PTM packages, expecting it to be positive (popular packages are fine-tuned).
5.4.1. Methods
We present the metrics we developed for claim 2. Metric 2: Turnover of Top PTMs: Drawing on prior work characterizing the stability of top packages over time (Dey and Mockus, 2018), we measured the top-$K$ turnover for each registry. Let $S_{current}$ be the set of $K$ most popular packages in the current snapshot, and $S_{last}$ be the set of top $K$ packages in the last snapshot. We also consider $S_{history}$, the set of packages in any snapshot before $S_{last}$. We then distinguish three categories for packages in $S_{current}$:
  1. (1)
    Remained: Packages that were in both $S_{last}$ and $S_{current}$.
    (2) $\text{Remained}=S_{last}\cap S_{current}$
  2. (2)
    Newcomers: Packages in $S_{current}$ in no previous snapshot.
    (3) $\text{Newcomers}=S_{current}\setminus(S_{history}\cup S_{last})$
  3. (3)
    Returning: Packages that were once in the top 1000 ($S_{history}$), were not in $S_{last}$, but are in $S_{current}$ again.
    (4) $\text{Returning}=(S_{history}\setminus S_{last})\cap S_{current}$
For the measurement, we defined popularity by the number of downloads, and examined the top-1000 packages. We obtained snapshots across four dates for both traditional software and PTMs. Data came from several HuggingFace snapshots (datasets: HF Model Metadata–June 2022, PTMTorrent–May 2023, PeaTMOSS–Oct. 2023). We took a current snapshot of the top 1000 PTMs directly from Hugging Face (April 2024). We used the Ecosyste.ms dataset (NPM, PyPI) for a comparison with traditional software package registries.
Metric 3: Number of Descendents of Top Packages: Our second measure of the impact of popularity on reuse was the number of descendent models. In this case, we defined descendent models as a downstream model that is fine-tuned and references the original model as a base model. We compare the number of descendent models with the popularity of model and determine the strength of correlation between them — the claim implies it should be positive.
For the measurement, we again defined popularity by the number of downloads. The descendent-base relation is available from the PeaTMOSS dataset, for the 15,000 most popular PTMs on Hugging Face. Given that these models account for $\sim$99of the downloads in the snapshot, we believe this is representative.
We initially planned to compare our findings with traditional software, similar to Metric 2. However, identifying a direct counterpart to PTM descendants in traditional software proved challenging. We considered using GitHub forks and GitHub or registry dependencies as analogs, but each presented unique implications that complicated a direct comparison with PTM descendants.
5.4.2. Results
Metric 2: Turnover of Traditional Software and PTM registries: Figure 6 shows the results for Hugging Face. The Hugging Face data does not match our interpretation of the claim. About half of the top-1K Hugging Face PTMs turned over in each snapshot. This high turnover rate suggests that packages on Hugging Face have a shorter lifespan, indicating a dynamic PTM environment where the requirements and preferred models frequently change (Gopalakrishna et al., 2023). Further investigation could analyze the lifecycle of these PTMs to determine whether they are newer models briefly appearing or established ones losing prominence, as well as identifying the traits of PTMs or maintainers who consistently stay within the top rankings. This result is consistent with the claim, but implies that popularity is likely driven by performance (a latent variable not present in the claim).
Figure 6 also shows NPM for comparison. The results for PyPI were similar. Traditional registries show stability with their most popular packages. Quantitatively, 2535 distinct packages were in the HuggingFace top-1K, while only 1127 were there for NPM.
Figure 6. The turnover of the top-1000 Packages on Hugging Face, NPM. Note that in traditional software package registries such as NPM, the level of turnover is low. Hugging Face has a much larger amount of packages re-enter the Top 1000 as shown by the larger green bars in the second and fourth snapshot compared to NPM, with few returners as shown by the relative lack of green bars.
The rapid turnover on Hugging Face compared to the stability observed in traditional software packages indicates that PTM needs are continually evolving, unlike the more established needs in traditional software domains. This evolution often leads to older PTMs being supplanted by newer models that better meet current requirements or offer superior performance, highlighting a market driven by innovation and rapid adaptation. This trend suggests that PTM users are quick to adopt new advancements, reflecting the field’s fast-paced development and the shifting demands across its various application domains.
Metric 3: Number of Descendents of Top Packages: Figure 7 shows the results for this metric. We observe a weak positive correlation between popularity (downloads) and the number of descendent models. Popular models tend to have more descendants, but the correlation was weaker than expected. Notably, some highly popular models had relatively few descendants, while others with similar popularity levels had many. This suggests that while popularity influences the likelihood of a model being chosen for further development, it is not the sole factor. Further analysis, including a breakdown by model task and domain, is necessary to understand if variations in descendant counts are influenced by the specific popularity within less prominent domains or tasks. This analysis could expose the decision-making processes engineers use when selecting models to fine-tune and develop further.
Figure 7. A heatmap showing the correlation between Downloads and Number of Direct Descendent Models. Note that this is a log-log plot. The red-line displays the best-fit relationship between downloads and descendant count, with a slope of $0.15$. This positive correlation suggests that models with higher download counts generally have a larger number of direct descendants, indicating that popular models tend to be reused more frequently in derivative work.
5.5. $C_{3}$: Docs Quality Impacts Model Selection
Our interpretation of Claim 3 is that PTMs with better documentation will be more popular.
5.5.1. Method
Metric 4: Documentation Quality: Prior work has examined documentation quality in many ways (Gong et al., 2023b; Tang and Nadi, 2023). We used those ideas to develop our measure of quality. The primary documentation for PTMs is called a “model card”, which is similar to the README of a GitHub repository or the landing page of an NPM or PyPI package. We considered two factors: (1) the completeness of the model card; and (2) the availability of metadata.
To measure completeness, we identified five typical sections found in highly popular PTMs such as Google’s Bet base uncased: Model Description, Limitations, How to Use, Training, and Evaluation. We scored model cards on an integer scale from 0 to 5 — to receive a score of 5, a PTM’s card needed all of these sections. To assess whether each section was present, we queried OpenAI’s ChatGPT-4 with the prompt shown in Listing LABEL:lst:prompt4docquality.
Listing 1: ChatGPT-4 prompt for evaluating model cards.
You will receive a model card and are expected to analyze it for the following details:
1. Model description: A description of the model itself
2. Limitations: Any limitations of the model
3. How to use: Instructions on how to use the model downstream
4. Training: Details of the training process or data
5. Evaluation: Reports on the models performance evaluation
Please respond with a JSON object indicating whether each of these points is present with true/false.
Here is the model card to evaluate:
To measure the availability of metadata, we referenced the PeaTMOSS dataset, which extracted over 20 distinct pieces of metadata if they were present in a PTM’s model card and associated configuration files. We scored PTMs on an integer scale from 0 to the maximum, i.e., the number of distinct pieces of metadata considered in the PeaTMOSS database schema. A PTM scoring perfectly in this category would possess all available metadata according to PeatMOSS.
The PeaTMOSS dataset comprises extracted metadata from all models on Hugging Face and includes snapshots of the top 15,000 most popular models, each with over 50 monthly downloads. These models were further analyzed using an LLM, which extracted additional metadata from the model cards and the config.json files. Consequently, our analysis focused on these top 15,000 models to assess whether documentation quality influences model selection.
For this measure, we considered popularity in three ways: Downloads and Likes, according to Hugging Face, and Downstream Dependents, based on the mapping offered by the PeaTMOSS dataset.
To test our interpretation of the metric, we selected the top 1000 and bottom 1000 models from PeaTMOSS’s set of 15,000 PTMs. We evaluated their documentation quality, summing an overall documentation score as a (0,1) metric that normalized and then weighted the two components equally. Then we compared the distributions using a box-and-whisker plot and statistical tests.
5.5.2. Results
Metric 4: Documentation Quality: We evaluated the impact of documentation quality on the popularity of PTMs. Two representative box and whisker plots are shown in Figure 8. Specifically, the top 1000 most popular models consistently exhibited significantly better documentation than the bottom 1000 models ($p<0.01$. This finding supports the claim that documentation quality significantly influences model selection. If the relation is causative (documentation $\rightarrow$ popularity), then research can focus on developing tools and methods to enhance the documentation quality of models, supporting better usability and adoption.
Figure 8. This figure shows the impact of documentation quality on model popularity using two popularity metrics: downloads and downstream reuse. The left box plot compares the documentation quality of the top 1,000 and bottom 1,000 models based on the number of downloads. The right box plot makes a similar comparison based on the number of downstream repositories. In both metrics, the top models demonstrate higher documentation quality scores than the bottom models, highlighting that models with better documentation are more popular and are reused more frequently.
6. Threats to Validity
We discuss three types of threats to validity (Wohlin et al., 2012), while considering the criticisms of Verdecchia et al. (Verdecchia et al., 2023).
Construct Threats are potential limitations of how we operationalized concepts. In the systematic literature review (RQ1), we manually extracted claims from papers, which might introduce potential bias to our results. As a mitigation, to improve objectivity two authors worked together on the process and the filtering of claims. In the validation of non-quantified claims, we proposed metrics that seemed suitable based on our judgment. As a mitigation, where possible we used multiple measures and leveraged previously-defined metrics.
Internal threats are those that affect cause-effect relationships. We emphasize that our approach for RQ2 is of the form: “If claim $C$ is true, then measurement $M$ should show us that…”. In each case the measurement produced the expected result. However, this result is correlative, not causative — there may be a latent variable in each case, or the causative relationship may be reversed. For example, in Figure 8 we found that better documentation correlates with greater popularity. It may be that the latent variable is performance, such that models become popular because they have good performance, and they accrue documentation because they are popular. When qualitative and quantitative claims agree, as is the case in this study, we learn both “Why?” and “How much?”.
External threats may impact generalizability. We recognize both immediate and longitudinal threats in this regard. Immediately, we were interested in studying PTM reuse, but we only examined one registry for PTMs: the Hugging Face platform. While this is by far the most popular and feature-rich platform, other platforms exist, such as PyTorch Hub (less popular), PapersWithCode (fewer features), and GitHub (not PTM-specific). In terms of longitudinally, keep in mind that Hugging Face is itself relatively young — created in 2016 and only seeing major use beginning in 2020 — so developer practices may not have stabilized. Lastly, we consider that the technologies and platforms that support PTMs are rapidly evolving, so current claims (whether qualitative or quantitative) may change over time and require ongoing reassessment. As indicated in the discussion, this property suggests an opportunity for further research, but it also means that our findings may be unstable.
7. Future Work
In Table 4, we identified one claim that we could not operationalize for measurement: $C_{4}$, that different deep learning-specific attributes would affect PTM selection and reuse. Although the PeaTMOSS dataset does include some of these measures in a structured format, the claim is sufficiently broad that quantifying it was beyond the scope of this study. Future work could explore the multi-variable relationship implied by this claim.
Our study provides the first systematic literature review of current knowledge about PTM reuse. Another opportunity is a systematic comparison of the software package registries associated with traditional software packages, and the software package registries associated with PTMs. Jiang et al. observed similarities and differences in the reuse processes (Jiang et al., 2023b) — how and to what extent can we measure the differences? As discussed in section 5.4.1 with respect to an analogue for PTM descendants, finding the limits of comparison (e.g., appropriate measurements) between traditional vs. PTM software package registries is an open challenge.
The rapid development of PTM technologies presents an ongoing challenge for empirical research on PTM reuse. The Hugging Face platform continues to grow exponentially (Castaño et al., 2023), the state of the art performance of models at all sizes continues to advance (pap, [n. d.]), and the tooling available to adapt and deploy these models continues to improve (Jajal et al., 2023). While datasets and tooling such as HFCommunity and PeaTMOSS support studies like ours, they also present some limitations. Innovation is needed to help empirical software engineering researchers keep up with the scale of activity and volume of data that we are seeing in the context of PTM reuse. Given Hugging Face’s dynamic growth, even data that is a few months old may not accurately reflect the current state of the platform, suggesting a need for regular snapshotting, which imposes significant storage requirements (beyond the already-substantial requirements of $>$50 TB). This highlights the need for tools that can provide real-time, incrementally-updated data to keep pace with rapid changes, ensuring that analyses remain relevant and reflective of the present situation in PTM reuse.
8. Conclusion
Pre-trained models are the motive force of the new generation of software engineering. Understanding engineers’ PTM reuse practices is crucial to optimizing and securing the process. Our systematic literature review and quantification of claims has illuminated significant aspects of PTM reuse. We also shed light on unique dynamics within the PTM reuse landscape as compared to traditional software package registries. Specifically, we observed a shorter lifespan of packages in PTM registries compared to traditional SPRs. Our findings underscore the need for research infrastructure and novel tools to support PTM reuse, adapting to the much higher turnover of popular PTM packages. We must ensure that PTM registries can meet the evolving demands of the software engineering community.
9. Data Availability
An anonymous artifact containing the results of the systematic literature review (RQ1) as well as our software and results for quantification of claims (RQ2) is available at https://github.com/anonsub1234/ptm-quantify-esem-2024.
10. Research Ethics
No human subjects were involved in the conduct of this project. We are aware of no other ethical concerns.
Acknowledgments
We acknowledge support from NSF awards OAC-2107020 and OAC-2107230.